OpenStudy (1dfanatic):

A jet leaves the Charlotte, North Carolina, airport traveling at an average rate of 564 km/h. Another jet leaves the airport one half hour later traveling at 744 km/h in the same direction. Write an equation to find how long the second jet will take to overtake the first. Solve your equation. Make sure to show your work for credit.

OpenStudy (danjs):

distance = rate*time

OpenStudy (1dfanatic):

I have to fill in a chart and it's confusing me lol

OpenStudy (danjs):

D = 564*t d = 744*t

OpenStudy (danjs):

If you think about it, if the first jet is traveling at that rate, then in 1/2 hour he will be 564/2 = 282 km distance. At the same time the second jet will take off... needs to gain 282 km to have the distances for both planes to be the same